OpenStudy (anonymous):

What is the graph of the absolute value equation? y = –1/2 |x|

jimthompson5910 (jim_thompson5910):

take the absolute value graph y = |x| and compress it by a factor of 2 to get y = 1/2|x| then flip it over the x-axis to get y = -1/2|x|

OpenStudy (anonymous):

So is it this one?

jimthompson5910 (jim_thompson5910):

it would have to go through (2, -1), so make sure it does

OpenStudy (anonymous):

Could you show me what you mean?

jimthompson5910 (jim_thompson5910):

when x = 2, y is -1 because y = -1/2|x| y = (-1/2)*|2| y = (-1/2)*2 y = -1

jimthompson5910 (jim_thompson5910):

so that's why the graph goes through the point (2, -1)